OpenStudy (anonymous):

What annual rate of interest is required to triple an investment in 11 years ?

OpenStudy (anonymous):

Do you know the formula for compounded interest?

OpenStudy (anonymous):

@leibystrauss umm I have written down 3p=p (1+r/1) ^11

OpenStudy (anonymous):

Have you been using \[A = \left( 1+\frac{ r }{ n } \right)^n\]

OpenStudy (anonymous):

Yes @leibystrauss

OpenStudy (anonymous):

Ok. Does the equation you wrote earlier make sense?

OpenStudy (anonymous):

@leibystrauss I don't know because I'm following my notes I took in class

OpenStudy (anonymous):

This formula is not for compounded interest that is not continuous; in the formula that you wrote interest would be added once a year.

OpenStudy (anonymous):

But it says what annual rate @leibystrauss

OpenStudy (anonymous):

I think the formula needs to be re-written \[A = \left( 1 + \frac{ r }{ n } \right)^{nt}\] P = principal or initial investment If the money triples it is 3 p t = time in years r = rate the n can be though of as periods r/1 means there is one period a year, the rate is applied once a year So essentially r/1 = r Since there is one period in the year nt = 1(t) = t t = 11 because it's 11 years Now you need to solve for r Does it make sense now?

OpenStudy (anonymous):

Yes, would I divide by three or get rid of the exponent first? @leibystrauss

OpenStudy (anonymous):

r/1 = r \[3p = p \left( 1 + r \right)^{11}\] what can you divide both sides by?

OpenStudy (anonymous):

P? @leibystrauss

OpenStudy (anonymous):

Correct, so re-write it.

OpenStudy (anonymous):

@leibystrauss 3=(1+r) ^11

OpenStudy (anonymous):

Ok. So how can you get rid of the 11th power?

OpenStudy (anonymous):

If \[8 = x^3\] How would you solve for x?

OpenStudy (anonymous):

Don't you cube root?

OpenStudy (anonymous):

You are technically correct. But I was thinking of something else. If \[16 = x^4\] How would you solve for x using a calculator?

OpenStudy (anonymous):

I'm not sure

OpenStudy (anonymous):

To remove the power of 4, you can raise both sides to the inverse of 4. Meaning raise both sides to 1/4 \[16^{\frac{ 1 }{ 4 }} = (x^{4})^{\frac{ 1 }{ 4 }}\] \[=> 16^{\frac{ 1 }{ 4 }} = x^1\]

OpenStudy (anonymous):

Ohh I see

OpenStudy (anonymous):

So what is the next step in your problem?

OpenStudy (anonymous):

What that mean I'd do 3^1/11=(1+r) ?

OpenStudy (anonymous):

Correct. And the final step?

OpenStudy (anonymous):

Subtract 1 from both sides ?

OpenStudy (anonymous):

Yes. You technically have your answer now. If you need the rate in % form then multiply your final answer by 100. What did you get as your final answer?

OpenStudy (anonymous):

Um first I got a negative number so that seemed wrong, so I tried it different and then I got 0.1=r

OpenStudy (anonymous):

What do you have after the 1? Recall 6.4% = 0.064 so you can't round 0.064 to 0.1

OpenStudy (anonymous):

I did 3^(1/11) && got 1.10 then I subtracted 1 && got 0.1

OpenStudy (anonymous):

You should have another digit after the 0

OpenStudy (anonymous):

1.105

OpenStudy (anonymous):

Correct, so if you do multiply it by 100 rate = 10.5%

OpenStudy (anonymous):

Thank you so much, I appreciate it (: